OpenStudy (anonymous):

how to find the area of a rhombus with vertices: (-1,-1) , (9,4) , (20,6) and (10,1)

OpenStudy (lgbasallote):

first step is to graph it...

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

but I don't know what to do next

OpenStudy (lgbasallote):

can you draw a rough sketch?

OpenStudy (lgbasallote):

here

OpenStudy (lgbasallote):

can you label the points too?

OpenStudy (anonymous):



OpenStudy (anonymous):

now use the distance formula to find the distance between (9,4) and (10,1) then you will have two nice easy triangles you can easily use A=1/2bh to solve for. then multiply by 2 to find the whole shape's area

OpenStudy (anonymous):

Do you know cross product?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

It is easier with cross product. You have to add 0 for z for each point. So you will be in 3D. \[ Area=\frac 1 2\left| AB \times AC \right| \]

OpenStudy (anonymous):

i thought it was the area is just the magnitude of the cross product, not half of it.... but i took vector calculus some yrs ago...

OpenStudy (anonymous):

this is for pre-calculus

OpenStudy (anonymous):

You can take A={-1, -1, 0} B={9, 4, 0} C={20, 6, 0}

OpenStudy (anonymous):

You are right. I was think about the triangle.

OpenStudy (anonymous):

indeed.

OpenStudy (anonymous):

\[ Area=\left| AB \times AC \right| \]

OpenStudy (anonymous):

I am not saying that one has to do it this way.

OpenStudy (anonymous):

how would you plug the numbers in?

OpenStudy (anonymous):

For precalculus, use the distance formula.

OpenStudy (anonymous):

I used it to split the rhombus in half to make 2 triangles

jimthompson5910 (jim_thompson5910):

Alternatively, you can use the formula A = (pq)/2 where p and q are the lengths of the diagonals. So you'll need to use the distance formula to find the lengths of the diagonals.